I would appreciate it if you can help me with this problem.
A water tank has an inlet pipe and a drain pipe. A full tank can be emptied in 30 min if the drain is opened and an empty tank can be filled in 45 min. withthe inlet pipe opened. If both pipes are accidentally opened when the tank is full, then how long will it take to empty the tank?
Do I add 30 min with 45 min?

Answer by mananth(16949)

A water tank has an inlet pipe and a drain pipe. A full tank can be emptied in 30 min if the drain is opened and an empty tank can be filled in 45 min. withthe inlet pipe opened. If both pipes are accidentally opened when the tank is full, then how long will it take to empty the tank?

Consider filling or emptying the complete tank as 1 job
1 job of emptying in 30 minutes
in 1 hour it does 2 jobs of emptying
1 job of filling it does in 45 minutes = 3/4 hours
in 1 hour it will do 4/3 jobs of filling
when both are opened 2-4/3 jobs of emptying done in 1 hour
=6-4/3
=2/3 job of emptying in 1 hour
1 job will be done in 3/2 hours
The tank will get emptied in 3/2 hours